# Instruction-tuning Aligns LLMs to the Human Brain

### v1

@@ Appendix A Language Models: Parameter count and Number of Layers
Table 3: Parameter count and number of layers for all 25 vanilla and instruction-tuned LLMs. The upper part contains encoder-decoder models of the T5 family, the lower parts decoder-only models of the LLaMA family. For the parameter count, “M” refers to million and “B” refers to billion. The number of layers for T5 models is a sum of the number of encoder and decoder layers.
@@ Appendix B Language Models: Links to models weights
Table 4: Link to model weights for all 25 vanilla and instruction-tuned LLMs. The upper part contains encoder-decoder models of the T5 family, the lower parts decoder-only models of the LLaMA family. We provide these links for reproducibility purposes.
@@
Appendix C MMLU and BBH benchmarks
@@
Appendix D Code Repositories
@@
We use Instruct-Eval repository to evaluate MMLU and BBH scores. Link: www.github.com/declare-lab/instruct-eval.
We use Stanford Alpaca repository for instruction-tuning. Link: www.github.com/tatsu-lab/stanford_alpaca).
Appendix E Results for Brain alignment
Table 5: Brain alignment results for all 25 vanilla and instruction-tuned LLMs. We provide these results for reproducibility purposes.
@@ Appendix E Results for Brain alignment
Table 6: Noise ceiling estimates for all 3 neural datasets. For Pereira2018 and Blank2014, noise ceiling estimates are computed using the Brain-Score repository, with details provided in Schrimpf et al. (2021). For Wehbe2014, noise ceiling estimates are also computed using a similar procedure.
@@
Appendix F Results for Next-word prediction, MMLU, BBH
Table 7: WikiText-2 NWP loss, MMLU Overall Score, and BBH Overall Score for all instruction-tuned LLMs. Results for vanilla LLMs are not shown as they are not adapted for the question formats in the MMLU and BBH benchmarks. We provide these results for reproducibility purposes.
WikiText-2 NWP Loss MMLU Overall Score BBH Overall Score
@@ Appendix F Results for Next-word prediction, MMLU, BBH
Notes on comparing next-word prediction (NWP) loss across model families.   The T5 and LLaMA models belong to separate model families. We wish to caution that comparing next-word prediction loss across different model families may not be meaningful. This is due to several reasons related to architectural differences, training methodologies, and objectives. (1) Architecture: T5 models have an encoder-decoder architecture while LLaMA models have a decoder-only architecture. (2) Training Objectives: The T5 models were trained on supervised and unsupervised tasks, while the LLaMA models were trained only on unsupervised text 3. (3) Loss computation: The loss functions for both model families are computed differently, making it inappropriate to directly compare their loss values. (4) Evaluation Metrics: Next-word prediction loss is just one metric, and it may not capture the overall language understanding capabilities of a model. Hence, we additionally evaluate these LLMs’ alignment to human brain activity, as well as their performance on problem-solving abilities (BBH) and tasks requiring world knowledge (MMLU). In summary, while NWP loss is a valuable metric for evaluating language models within the same family or architecture, comparing across different model families may not be meaningful.
Appendix G Results for Correlations of Brain Alignment with LLM properties
@@
Figure 4: Correlation between brain alignment and various LLM properties: (A) MMLU benchmark global score, (B) BBH benchmark score with only world knowledge tasks, (C) number of parameters of the model, and (D) Next word prediction (NWP) performance.
@@
Figure 5: Correlation between brain alignment and various LLM properties for all 25 LLMs: (A) number of parameters of the model, and (B) Next word prediction (NWP) performance.
Appendix H Results for Behavioral alignment
Table 8: Behavioral alignment results for all 25 vanilla and instruction-tuned LLMs. We provide these results for reproducibility purposes.
@@ Appendix H Results for Behavioral alignment
Appendix I Results for Instruction-tuning LLaMA-7B on Alpaca dataset
Instruction model
We instruction-tune LLaMA-7B on the Stanford Alpaca dataset (Taori et al., 2023) using the default training process, following the code in www.github.com/tatsu-lab/stanford_alpaca. In particular, the model is instruction-tuned using 5252K instruction-following examples generated through methods inspired by Self-Instruct (Wang et al., 2022a)). This model is labeled “Instruction” in Figure 6.
No instruction model (Ablation)
We also train an ablation model with the same process and training data as the default instruction-tuning, but remove the instruction portion from each training sample. This ablation model is labeled “No instruction” in Figure 6. This ablation experiment disentangles: (1) training data (present in both Instruction and No instruction), from (2) training LLMs to understand and follow instructions (present only in Instruction).
We use all provided training samples from the Alpaca dataset, thus ensuring that the models are trained on the same amount of data. We observe that brain alignment of the “No Instruction” ablation model increases during fine-tuning but stays lower than its “Instruction” counterpart.

### v2

@@ Appendix A Language Models: Parameter count and Number of Layers
gpt2-small 124 M 12
gpt2-small-alpaca 124 M 12
gpt2-medium 355 M 24
gpt2-medium-alpaca 355 M 24
gpt2-large 774 M 36
gpt2-large-alpaca 774 M 36
gpt2-xl 1.5 B 48
gpt2-xl-alpaca 1.5 B 48
Table 3: Parameter count and number of layers for all vanilla and instruction-tuned LLMs. For the parameter count, “M” refers to million and “B” refers to billion. The number of layers for T5 models is a sum of the number of encoder and decoder layers.
@@ Appendix B Language Models: Links to models weights
gpt2-small https://huggingface.co/openai-community/gpt2
gpt2-small-alpaca https://huggingface.co/vicgalle/gpt2-alpaca
gpt2-medium https://huggingface.co/openai-community/gpt2-medium
gpt2-medium-alpaca https://huggingface.co/linkanjarad/GPT2-Medium-Alpaca-355m
gpt2-large https://huggingface.co/openai-community/gpt2-large
gpt2-large-alpaca https://huggingface.co/reasonwang/gpt2-large-alpaca
gpt2-xl https://huggingface.co/openai-community/gpt2-xl
gpt2-xl-alpaca https://huggingface.co/Rachneet/gpt2-xl-alpaca
Table 4: Link to model weights for all vanilla and instruction-tuned LLMs. We provide these links for reproducibility purposes.
Appendix C Method for computing Brain alignment
Figure 4: Method for computing Brain alignment, the similarity of an LLM’s internal representations to human brain activity.
We follow a general approach for aligning LLM representations to fMRI brain activity used in prior works (Jain & Huth, 2018; Toneva & Wehbe, 2019; Schrimpf et al., 2021). The approaches used for the three neural datasets are similar. Hence, we explain the main details only for Wehbe2014, followed by key differences for the other two datasets.
C.1 Wehbe2014
Wehbe2014 (Wehbe et al., 2014): The data includes fMRI recordings of 8 human participants reading chapter 9 of the book Harry Potter and the Sorceror’s Stone (Rowling et al., 1998). Participants read the chapter at a fixed interval of one word every 0.5 seconds. The chapter contains 5,176 words.
LLM representations. We extract LLM representations of the text at each layer $\in\mathbb{R}^{w\times d}$, where $w$ is the number of words in the text and $d$ is the embedding size of the LLM.
fMRI activity. Data analyses were performed on fMRI BOLD signals extracted from the language network for all three neural datasets. fMRI activity was recorded at regular intervals of 2 seconds, with a total of 1,351 recorded time-points (TRs, times of repetition). We construct a stimulus-response matrix for fMRI activity $\in\mathbb{R}^{TR\times v_{i}}$, where $TR$ is the number of TRs and $v_{i}$ is the number of voxels in the brain of participant $i$.
Alignment of LLM representations to fMRI activity. We use PCA to reduce the dimensionality of the LLM representations to result in a matrix $\in\mathbb{R}^{w\times 10}$. As the number of words in the text is greater than the number of TRs, we down-sample the LLM word-level representations to the TR rate by averaging the LLM representations to the corresponding TRs, producing a matrix $\in\mathbb{R}^{TR\times 10}$. As the response recorded by fMRI peaks about 6 seconds after stimulus onset, we follow prior methods by including preceding time-points for each time-point. The final LLM representations are constructed by concatenating, for each TR, the LLM representations corresponding to the previous 4 TRs, producing a matrix $\in\mathbb{R}^{TR\times 40}$. This accounts for the lag in the hemodynamic response that fMRI records. Finally, we learn a linear function, regularized by the ridge penalty, that uses the modified LLM representations $\in\mathbb{R}^{TR\times 40}$ to predict the fMRI activity of participants $\in\mathbb{R}^{TR\times v_{i}}$. We train the function using 4-fold cross-validation, where each fold corresponds to a separate run of fMRI collection. We also remove 10 TRs from the LLM representations and fMRI activity from the ends of each test run to avoid train-test overlap. To compute the brain alignment, we evaluate the Pearson correlation (r) between the predicted fMRI activity $\in\mathbb{R}^{TR\times v_{i}}$ and recorded fMRI activity $\in\mathbb{R}^{TR\times v_{i}}$.
C.2 Blank2014
Blank2014 (Blank et al., 2014): The data consists of fMRI recordings of 5 human participants listening to 8 naturalistic stories from the Natural Stories Corpus (Futrell et al., 2018).
fMRI activity was recorded at regular intervals of 2 seconds, with a total of 1,317 TRs. We average the BOLD signals across voxels within each language-responsive region of interest (ROI) of each participant to increase the signal-to-noise ratio, following Schrimpf et al. (2021). As there are 60 ROIs in total across the 5 participants, this produces a stimulus-response matrix $\in\mathbb{R}^{1,317\times 60}$.
C.3 Pereira2018
Pereira2018 (experiments 2 and 3 from Pereira et al., 2018): In experiment 2, 9 participants read 384 sentences taken from 96 text passages. In experiment 3, 6 participants read 243 sentences from 72 text passages. Each sentence was displayed for 4 seconds on a screen.
We average the fMRI responses for each sentence, resulting in one data point per sentence per language-responsive voxel of each participant. For experiment 2, there are 384 sentences and 12,195 language-responsive voxels across the participants. For experiment 3, there are 243 sentences and 8,121 language-responsive voxels across the participants. We concatenate the fMRI responses across sentences and participants for each experiment. This produces a stimulus-response matrix $\in\mathbb{R}^{384\times 12,195}$ for experiment 2 and a stimulus-response matrix $\in\mathbb{R}^{243\times 8121}$ for experiment 3.
Appendix D Method for computing Behavioral alignment
Figure 5: Method for computing Behavioral alignment. The same language stimuli are presented to LLMs and human participants, using the Futrell2018 benchmark in Brain-Score, which contains naturalistic stories. We compute the behavioral alignment as the Pearson correlation between LLM perplexity for each word and human reading times for each word.
Futrell2018 (Futrell et al., 2018): The data consists of reading times (RTs) of 179 participants on a total of 10,256 words across 485 sentences, taken from 10 stories from the Natural Stories Corpus (Futrell et al., 2018). The stories were presented online to Amazon Mechanical Turk users visually in a dashed moving window display. Participants press a button to reveal each consecutive word. When they press the button, the most recent word is converted back to dashes, and the next word is revealed. The time between button presses offers an estimate of comprehension difficulty (Ehrlich & Rayner, 1981; Hale, 2001; Smith & Levy, 2013). We construct a stimulus-response matrix of RTs $\in\mathbb{R}^{10,256\times 179}$.
We provide the same stimuli to an LLM and record its perplexity for each word, producing an LLM perplexity matrix $\in\mathbb{R}^{10,256}$. We compute behavioral alignment as the Pearson correlation between per-word LLM perplexity and human RTs. Perplexity for LLMs and reading times for humans offer insights into comprehension difficulty (Ehrlich & Rayner, 1981; Hale, 2001; Smith & Levy, 2013), allowing us to examine whether LLMs and humans share similarities in terms of which words and sentences they find challenging or surprising.
Appendix E MMLU and BBH benchmarks
@@
Appendix F Code Repositories
@@
We use the Instruct-Eval repository to evaluate MMLU and BBH scores. Link: www.github.com/declare-lab/instruct-eval.
We use the Stanford Alpaca repository for instruction-tuning. Link: www.github.com/tatsu-lab/stanford_alpaca).
Appendix G Results for Brain alignment
@@
gpt2-small 0.305 0.123 0.088 0.172
gpt2-small-alpaca 0.298 0.121 0.081 0.167
gpt2-medium 0.329 0.081 0.089 0.166
gpt2-medium-alpaca 0.325 0.080 0.090 0.165
gpt2-large 0.342 0.077 0.101 0.173
gpt2-large-alpaca 0.336 0.074 0.101 0.170
gpt2-xl 0.358 0.140 0.102 0.200
gpt2-xl-alpaca 0.343 0.139 0.093 0.192
Table 5: Brain alignment results for all vanilla and instruction-tuned LLMs. We provide these results for reproducibility purposes.
@@
Table 6: Noise ceiling estimates for all 3 neural datasets. fMRI measurements inherently include noise, i.e., fluctuations not due to neurons firing, from sources such as environmental factors, the scanner, and movements of human participants. Consequently, a “noise ceiling” is often computed by recording multiple samples from the same subjects for the same stimuli, providing an upper-limit estimate of the possible correlation in fMRI activity. In our study, we followed prior work for calculating the noise ceiling. For Pereira2018 and Blank2014, noise ceiling estimates are computed using the Brain-Score repository, with details provided in Schrimpf et al. (2021). For Wehbe2014, noise ceiling estimates are computed using a similar procedure.
Appendix H Results for Next-word prediction, MMLU, BBH
WikiText-2 NWP Loss MMLU Score BBH Score
@@ Appendix H Results for Next-word prediction, MMLU, BBH
gpt2-small-alpaca 4.193 0.270 0.273
gpt2-medium-alpaca 3.333 0.260 0.276
gpt2-large-alpaca 3.614 0.260 0.278
gpt2-xl-alpaca 3.283 0.270 0.284
Table 7: WikiText-2 NWP loss, MMLU Overall Score, and BBH Overall Score for all instruction-tuned LLMs. Results for vanilla LLMs are not shown as they are not adapted for the question formats in the MMLU and BBH benchmarks. Results in gray are close to random performance. We provide these results for reproducibility purposes.
@@
Notes on comparing next-word prediction (NWP) loss across model families.   The T5 and LLaMA models belong to separate model families. We wish to caution that comparing next-word prediction loss across different model families may not be meaningful. This is due to several reasons related to architectural differences, training methodologies, and objectives. (1) Architecture: T5 models have an encoder-decoder architecture while LLaMA models have a decoder-only architecture. (2) Training Objectives: The T5 models were trained on supervised and unsupervised tasks, while the LLaMA models were trained only on unsupervised text (Section 3). (3) Loss computation: The loss functions for both model families are computed differently, making it inappropriate to directly compare their loss values. (4) Evaluation Metrics: Next-word prediction loss is just one metric, and it may not capture the overall language understanding capabilities of a model. Hence, we additionally evaluate these LLMs’ alignment to human brain activity, as well as their performance on problem-solving abilities (BBH) and tasks requiring world knowledge (MMLU). In summary, while NWP loss is a valuable metric for evaluating language models within the same family or architecture, comparing across different model families may not be meaningful.
Appendix I Results for Correlations of Brain Alignment with LLM properties
@@
Figure 6: Correlation between brain alignment and various LLM properties: (A) MMLU benchmark overall score, (B) BBH benchmark score for world knowledge tasks, (C) number of parameters of the model, and (D) Next-word prediction (NWP) loss.
@@
Figure 7: Correlation between brain alignment and various LLM properties for all LLMs (including vanilla models): (A) number of parameters of the model, and (B) Next-word prediction (NWP) loss.
Appendix J Additional Brain Alignment results (Linear, RSA, CKA) for Gemma and LLaMA-2 models
Pretrained models
Model Linear CKA RSA
Gemma-2B 0.76 1.94 0.63
Gemma-7B 0.90 2.25 0.62
LLaMA-2-7B 0.97 1.81 0.65
LLaMA-2-13B 1.08 2.51 0.68
Instruction-tuned models
Model Linear CKA RSA
Gemma-2B 0.72 2.44 0.70
Gemma-7B 0.94 2.55 0.72
LLaMA-2-7B 1.02 2.00 0.63
LLaMA-2-13B 1.08 2.77 0.65
Table 8: Additional Brain Alignment results (Linear, RSA, CKA) for Gemma and LLaMA-2 models. “Linear” refers to the linear predictivity metric used in the main paper, normalized by the noise ceiling computed in Appendix G.
In Section 4, we study the representational similarity between LLMs and human brain activity (brain alignment) using the linear predictivity similarity metric (“Linear”). This metric involves training a linear regression model to predict fMRI brain activity based on LLM representations. Here, we additionally validate our results using the RSA and CKA similarity metrics (Table 8), confirming that the brain alignment results are consistent across different similarity metrics.
In Section 4.1, we observed that instruction-tuning generally improves brain alignment for models from the T5 and LLaMA-1 families. We further validate this trend with recently released models from the Gemma and LLaMA-2 families (Table 8). These results hold across various similarity metrics, including Linear predictivity, RSA, and CKA.
Appendix K Results for Instruction-tuning LLaMA-7B on Alpaca dataset
Instruction model.
We instruction-tune LLaMA-7B on the Stanford Alpaca dataset (Taori et al., 2023) using the default training process, following the code in www.github.com/tatsu-lab/stanford_alpaca. In particular, the model is instruction-tuned using $52$K instruction-following examples generated through methods inspired by Self-Instruct (Wang et al., 2022a). This model is labeled “Instruction” in Figure 8.
No instruction model (Ablation).
We also train an ablation model with the same process and training data as the default instruction-tuning, but remove the instruction portion from each training sample. This ablation model is labeled “No instruction” in Figure 8. This ablation experiment disentangles: (1) training data (present in both “Instruction” and “No instruction”), from (2) training LLMs to understand and follow instructions (present only in “Instruction”).
We use all provided training samples from the Alpaca dataset, thus ensuring that the models are trained on the same amount of data. We observe that brain alignment of the “No Instruction” ablation model increases during fine-tuning but stays lower than its “Instruction” counterpart. This shows that brain alignment improvements are due to both (1) training data (present in both models) and (2) the process of training LLMs to understand and follow instructions (present only in the “Instruction” model).
Appendix L Results for Behavioral alignment
@@
gpt2-small 0.367
gpt2-small-alpaca 0.335
gpt2-medium 0.350
gpt2-medium-alpaca 0.345
gpt2-large 0.331
gpt2-large-alpaca 0.338
gpt2-xl 0.318
gpt2-xl-alpaca 0.336
Table 9: Behavioral alignment results for all vanilla and instruction-tuned LLMs. We provide these results for reproducibility purposes.
@@
Appendix M Results for Correlations of Behavioral Alignment with LLM properties
Figure 9: Correlation between behavioral alignment and various LLM properties: (A) MMLU benchmark overall score, (B) number of parameters of the model, (C) Next-word prediction (NWP) loss for T5 models, and (D) NWP loss for LLaMA models.